### v1

# BWTP model for hadrons $p_{\rm T}$-spectra and elliptic flow $v_{2}$ in AA collisions at the LHC

Abstract
A generalization of the phenomenological BWTP model, recently proposed by the author for the hadrons transverse momentum ($p_{\textrm{T}}$) spectra measured at the LHC, is developed to also describe the hadrons $p_{\textrm{T}}$-differential elliptic flow coefficients $v_{2}$ in identical nuclei (AA) collisions of different centralities. The model describes well the available data on $p_{\textrm{T}}$-spectra and $v_{2}$ for any $p_{\textrm{T}}$ of various particles, from pions to charmonia, in Pb–Pb at $\sqrt{s_{\mathrm{NN}}}=$ 2.76 and 5.02 TeV, and in Xe–Xe at $\sqrt{s_{\mathrm{NN}}}=$ 5.44 TeV. Also, predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$ 5.36 TeV are given. While the model is mainly targeting the LHC energies, it also works at much lower RHIC energies.
pacs:
24.10.Pa, 13.85.Ni, 12.38.Mh, 25.75.-q
I Introduction
Recently, the BWTP model (BWTPM) was proposed, which describes quite accurately the midrapidity $p_{\textrm{T}}$-spectra of various hadrons in the whole $p_{\textrm{T}}$ range measured in pp and Pb–Pb collisions at the LHC [1]. These $p_{\textrm{T}}$-spectra are integrated over the hadron azimuthal angle $\varphi$. They depend strongly on the collision centrality and centre-of-mass energy per nucleon-nucleon pair $\sqrt{s_{\mathrm{NN}}}$. BWTPM has three terms including a standard blast-wave model (BWM) term [2] to account for the hydrodynamical radial flow and kinetic freeze-out effects, a term in the Tsallis (Ts) distribution form [3, 4], responsible for the contribution from the resonance decays, and a power-law (Po) term describing the QCD hard processes. As shown in Ref. [1], the model parameters depend generally on the collision system (pp or Pb–Pb), but some of them are independent of the $\sqrt{s_{\mathrm{NN}}}$ or the particle type or the collision centrality.
In the present paper, a generalization of the BWTPM is developed to describe, in addition to the $p_{\textrm{T}}$-spectra, also the midrapidity elliptic flow coefficients $v_{2}$ in identical nuclei (AA) collisions. The $v_{2}$ characterizes the azimuthal asymmetry of hadron momentum distribution and depends on hadron type, $p_{\textrm{T}}$, collision system, and centrality. The asymmetry is driven mainly by an almond shape (in the plane transverse to the collision axis) of the hadronic medium (quark-gluon plasma) created in non-central collisions. Due to this shape, the in-medium pressure gradient and radial flow velocity become anisotropical resulting in the anisotropy of the hadron momentum spectra [5]. Following to Ref. [6], we use the generalized BWM with transversally anisotropic freeze-out hyper-surface of elliptical shape, instead of the isotropic circular one considered in Refs. [1, 2]. For the hadronic medium radial flow velocity we assume similar azimuthal dependence as in Refs. [6, 7]. Also, the other two therms of BWTPM are modified by including a simple dependence on $\varphi$. Having redefined so the model, the hadron $v_{2}$ can be determined as the second coefficient in the Fourier series of its invariant yield $\textrm{d}^{3}N/(p_{\textrm{T}}\textrm{d}p_{\textrm{T}}\textrm{d}y\textrm{d}\varphi)$ [6, 7, 8, 9, 10]
$\thinspace v_{2}=\int_{0}^{2\pi}d\varphi\cos(2\varphi)\frac{\textrm{d}^{3}N}{\textrm{d}p_{\textrm{T}}\textrm{d}y\textrm{d}\varphi}~\bigg{/}\int_{0}^{2\pi}d\varphi\frac{\textrm{d}^{3}N}{\textrm{d}p_{\textrm{T}}\textrm{d}y\textrm{d}\varphi}~,$ (1)
where $y$ is the hadron rapidity. The $v_{2}$ contribution dominates in this series for non-central collisions.
Next objective of the present paper is to reduce strongly the number of the BWTPM parameters by parameterizing their dependence on the collision system, centrality, energy, and on the particle type and $p_{\textrm{T}}$ using simple formulae. As a measure of collision centrality, the corresponding charged-particle multiplicity density at midrapidity ($\eta$ is pseudo-rapidity)
$M_{\rm ch}\equiv\langle\textrm{d}N_{\rm ch}/\textrm{d}\eta\rangle_{\eta=0}$ (2)
is used which serves as a good scaling variable for the comparison of hadron production mechanisms in different collision systems (see, e.g., the ALICE review paper [11]). These parameterizations increase significantly the predictive power of the model. In addition, we introduce another simplification of the BWTPM. While in Ref. [1] it has two different functional forms for $p_{\textrm{T}}\leq 40$ GeV and $p_{\textrm{T}}>40$ GeV, here it has the same form for any $p_{\textrm{T}}$.
The values of the BWTPM parameters are obtained then from a simultaneous fit of the combined data on different hadron $p_{\textrm{T}}$-spectra and $v_{2}$ measured mostly by the ALICE experiment in Pb–Pb and Xe–Xe collisions at the LHC. In addition to the hadrons considered in Ref. [1] ($\pi^{\pm}$, $\pi^{0}$, $\mathrm{K}^{\pm}$, ${\mathrm{K}}^{0}_{\mathrm{S}}$, $\mathrm{p}$, $\mathrm{\overline{p}}$, ${\mathrm{K}}^{*0}$, $\overline{\mathrm{K}^{*0}\mkern-15.0mu}$ , $\phi$, $\Lambda$, $\overline{\Lambda}$, $\Xi^{-}$, $\overline{\Xi^{+}\mkern-12.0mu}$ , $\Omega^{-}$, $\overline{\Omega^{+}\mkern-12.0mu}$ , ${\mathrm{D}}^{0}$, $\mathrm{D}^{\pm}$, $\mathrm{J}/\psi$, and unidentified charged-particles), the $\eta$, $\rho^{0}$ mesons and $\Lambda_{c}^{+}$$\overline{\Lambda^{-}\mkern-12.0mu}_{c}$ baryons are included too. We assume that a particle and its antiparticle have the same $p_{\textrm{T}}$-spectrum and $v_{2}$. While in the LHC energy domain, which is the main target of our study, this assumption is perfectly valid, it could be violated at much lower energies. To test our model at such energies, we included in the simultaneous fit also some limited data on $p_{\textrm{T}}$-spectra and $v_{2}$ measured at the RHIC.
Note that the previous similar simultaneous fits of the LHC data on hadrons $p_{\textrm{T}}$-spectra and $v_{2}$ were done mainly at low $p_{\textrm{T}}$ ($p_{\textrm{T}}\lesssim 3$ GeV/c[7, 11, 12] using various generalizations of the BWM. Contrary, our model, due to its additional power-law terms, is able to describe much larger data set in the whole available $p_{\textrm{T}}$-region.
The paper is organized as follows: Sec. II gives details of the generalized BWTPM. Sec. III presents a parameterization for the $M_{\rm ch}$ data measured in various AA collisions. Sec. IV describes parameterizations for the model parameters dependence on the AA collision characteristics and the particle type and $p_{\textrm{T}}$. In Sec. V the model free parameters are determined from the simultaneous fit of all the used data. In Sec. VI the results and predictions are discussed. Conclusions are given in the last section.
II Description of BWTPM
Present model consists of three terms which generalize the similar terms of Ref. [1]. The first (BW) term corresponds to the generalized BWM [6] and two others (named Ts and Po as in Ref. [1]) include additionally the second Fourier coefficients $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$, respectively, to account for the elliptic flow. Thus, for a particle $\varphi$-dependent $p_{\textrm{T}}$-spectrum in AA collisions at midrapidity, where it can be considered as independent of the rapidity $y$, we propose the following model
$\frac{\textrm{d}^{3}N}{\textrm{d}p_{\textrm{T}}\textrm{d}y\textrm{d}\varphi}\equiv F_{\varphi}\,$, (3)
$F_{\varphi}(p_{\textrm{T}})=\frac{g}{2\pi^{3}}\frac{f_{N}}{f_{S}}p_{\textrm{T}}m_{\textrm{T}}\left[f_{BW}V\int_{0}^{1}\hat{r}d\hat{r}\int_{0}^{2\pi}\frac{d\phi}{2\pi}\right$.
$\left.K_{1}(\gamma_{\textrm{T}}\frac{m_{\textrm{T}}}{T})\exp{(\gamma_{\textrm{T}}\beta_{\textrm{T}}\frac{p_{\textrm{T}}}{T}\cos(\varphi-\phi_{b}))}\right$.
$+\left.f_{1}V_{1}[1+2v_{2t}(p_{\textrm{T}})\cos(2\varphi)](1+\frac{m_{\textrm{T}}}{c_{1}e_{1}n})^{-c_{1}n}\right$.
$+\left.f_{2}V_{2}[1+2v_{2p}(p_{\textrm{T}})\cos(2\varphi)](c_{2}^{2}+\frac{p_{\textrm{T}}^{2}}{e_{2}^{2}})^{-\frac{n}{2}}\right],$ (4)
where N is the particle yield per collision event, $g=2J+1$ is its spin degeneracy factor, $m_{\textrm{T}}=\sqrt{m^{2}+p_{\textrm{T}}^{2}}$ is its transverse mass, $\gamma_{\textrm{T}}=1/\sqrt{1-\beta_{\textrm{T}}^{2}}$, $\beta_{\textrm{T}}$ is the hadronic medium radial (transverse) flow velocity at the kinetic freeze-out
$\beta_{\textrm{T}}=\tanh{\rho},\quad\rho=[\rho_{0}+\rho_{2}\cos(2\phi_{b})]\hat{r}^{k},$ (5)
$\rho$ is the flow rapidity with the second order anisotropy $\rho_{2}$ and profile exponent $k$, $V=\pi R_{x}R_{y}\tau_{f}$ is the volume of this medium with transverse sizes $R_{x}$ and $R_{y}$, $\tau_{f}$ and $T$ are its proper life-time and temperature, $K_{1}$ is a modified Bessel function of the second kind. Note that $\tanh{\rho_{0}}$ is equal to $\beta_{\textrm{s}}$ used in Ref. [1]. The global normalization parameter $f_{N}$ is close to unity at LHC energies but could depend on the normalization conditions of the used data obtained in different experiments. Its energy dependence is significant at lower RHIC energies and will be specified in Sec. IV. Parameter $f_{S}$ is equal unity (as in Ref [1]) for ordinary hadrons consisting of up and down quarks only. But for hadrons containing strange or heavier quarks it is greater unity and quantifies the suppression of their yields depending on the hadron type, $p_{\textrm{T}}$, and collision centrality. It describes the observed enhancement of the relative yields of such hadrons with respect to ordinary ones when going from peripheral to central collisions (i.e., from small to large $M_{\rm ch}$[11]. The explicit form of $f_{S}$ is specified in Sec. IV.
The azimuthal dependence of $\beta_{\textrm{T}}$ in Eq. (5) is similar to the one in Refs. [6, 7] while its dependence on radial variable $\hat{r}$ is defined for arbitrary $k$ [1, 2] which varies strongly with the collision centrality ($k=1$ in Refs. [6, 7]). The meaning of azimuthal angles $\phi_{b}$ and $\phi$ and their interrelation
$\tan\phi_{b}=\sqrt{r_{xy}}\tan\phi,~~r_{xy}\equiv R_{x}^{2}/R_{y}^{2}$ (6)
are explained in Refs. [6, 7]. In our model, three parameters $R_{x},R_{y}$, and $\tau_{f}$ enter into two independent combinations via parameters $V$ and $r_{xy}\leq 1$.
The important quantities for the elliptic flow $v_{2}$ are $\rho_{2},r_{xy}$ [6, 7], $k$, and functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ in Eq. (II), though the $T$ and $\rho_{0}$ also influence the $p_{\textrm{T}}$-dependence of $v_{2}$ [6]. From hydrodynamic models and experimental data it follows an approximate linear relation between $v_{2}$ and the eccentricity $\varepsilon_{2}$ of the collision geometry [13, 14, 15, 16], which can be defined in the present model as
$\varepsilon_{2}=(1-r_{xy})/(1+r_{xy}).$ (7)
Moreover, in the generalized BWM [6] with $\rho_{2}=0$ one gets $v_{2}=\varepsilon_{2}/2$. We have checked by fitting the $v_{2}$ data that parameter $\rho_{2}$ as well as the functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ have similar to $\varepsilon_{2}$ centrality dependence. For the model simplicity, we assume that they are proportional to $\varepsilon_{2}$. Their explicit forms will be specified in Sec. IV. All other parameters in Eq. (II) have the same meaning as in Ref. [1]. Note that parameters $f_{BW},f_{1},f_{2},c_{1}$, and $c_{2}$ are the same for any AA collision system and depend only on the particle type.
Now, integration of Eq. (3) by $\varphi$ gives
$\frac{\textrm{d}^{2}N}{\textrm{d}p_{\textrm{T}}\textrm{d}y}\equiv F\,$, (8)
$F(p_{\textrm{T}})=\frac{g}{\pi^{2}}\frac{f_{N}}{f_{S}}p_{\textrm{T}}m_{\textrm{T}}\left[f_{BW}V\int_{0}^{1}\hat{r}d\hat{r}\int_{0}^{2\pi}\frac{d\phi}{2\pi}\right$.
$\left.K_{1}(\gamma_{\textrm{T}}\frac{m_{\textrm{T}}}{T})I_{0}(\gamma_{\textrm{T}}\beta_{\textrm{T}}\frac{p_{\textrm{T}}}{T})+f_{1}V_{1}(1+\frac{m_{\textrm{T}}}{c_{1}e_{1}n})^{-c_{1}n}\right$.
$+\left.f_{2}V_{2}(c_{2}^{2}+\frac{p_{\textrm{T}}^{2}}{e_{2}^{2}})^{-\frac{n}{2}}\right].$ (9)
Performing similar integration in Eq. (1) one finds (see also Ref. [7])
$v_{2}(p_{\textrm{T}})=\frac{g}{\pi^{2}}\frac{f_{N}p_{\textrm{T}}m_{\textrm{T}}}{f_{S}F(p_{\textrm{T}})}\left[f_{BW}V\int_{0}^{1}\hat{r}d\hat{r}\int_{0}^{2\pi}\frac{d\phi}{2\pi}\cos(2\phi_{b})\right$.
$\left.K_{1}(\gamma_{\textrm{T}}\frac{m_{\textrm{T}}}{T})I_{2}(\gamma_{\textrm{T}}\beta_{\textrm{T}}\frac{p_{\textrm{T}}}{T})+f_{1}V_{1}v_{2t}(p_{\textrm{T}})(1+\frac{m_{\textrm{T}}}{c_{1}e_{1}n})^{-c_{1}n}\right$.
$+\left.f_{2}V_{2}v_{2p}(p_{\textrm{T}})(c_{2}^{2}+\frac{p_{\textrm{T}}^{2}}{e_{2}^{2}})^{-\frac{n}{2}}\right].$ (10)
The $I_{0}$ and $I_{2}$ in Eqs. (II) and (II) are modified Bessel functions of the first kind. So, we found Eqs. (8), (II), and (II) for hadrons $p_{\textrm{T}}$-spectrum and elliptic flow, respectively. They can be used to describe the corresponding data for different identified particles. There are also similar data sets for the unidentified charged particles which are a mixture of the $\pi^{\pm},\mathrm{K}^{\pm},\mathrm{p},\mathrm{\overline{p}}$, and charged hyperons. These data have usually much larger $p_{\textrm{T}}$ reach than the ones for the identified particles. Charged-particle $p_{\textrm{T}}$-spectra can be defined, assuming same spectrum for a particle and its anti-particle, by the equations [1]
$\frac{\textrm{d}^{3}N_{\rm ch}}{\textrm{d}p_{\textrm{T}}\textrm{d}\eta\textrm{d}\varphi}=2(\sum\limits_{i=\pi,\mathrm{K,p}}\frac{p_{\textrm{T}}}{{m_{\textrm{T}}}_{,i}}F_{\varphi,i}+f_{hyp}\frac{p_{\textrm{T}}}{{m_{\textrm{T}}}_{,\mathrm{p}}}F_{\varphi,\mathrm{p}})\,$, (11)
$\frac{\textrm{d}^{2}N_{\rm ch}}{\textrm{d}p_{\textrm{T}}\textrm{d}\eta}=2(\sum\limits_{i=\pi,\mathrm{K,p}}\frac{p_{\textrm{T}}}{{m_{\textrm{T}}}_{,i}}F_{i}+f_{hyp}\frac{p_{\textrm{T}}}{{m_{\textrm{T}}}_{,\mathrm{p}}}F_{\mathrm{p}})\,.$ (12)
Here, the functions $F_{\varphi,i}$ and $F_{i}$ are defined by Eqs. (II) and (II), respectively. The factors 2 account for the positive and negative particles and factors $p_{\textrm{T}}/m_{\textrm{T}}$ account for the change from the rapidity to pseudo-rapidity at midrapidity ($d\eta/dy\approx p_{\textrm{T}}/m_{\textrm{T}}$). The last terms in Eqs. (11) and (12) describe approximately the small contribution of hyperons via the proton contribution, scaled by parameter $f_{hyp}$. Using Eqs. (11) and (1) and performing integration by $\varphi$ one can find easily the analogue of Eq. (II) for the unidentified charged particles. Note that the Eqs. (8), (II) and (12) for the $p_{\textrm{T}}$-spectra are similar to the corresponding ones in Ref. [1] if $f_{N}=f_{S}=1$ and $\rho_{2}=0$.
We have checked that the presented generalized BWTPM is able to fit the LHC data on midrapidity hadron $p_{\textrm{T}}$-spectra and $v_{2}$ in Pb–Pb and Xe–Xe collisions with a good accuracy, similar to one achieved in Ref. [1]. However, as in Ref. [1], the number of the model free parameters is too big, which reduces the model’s predictive power. It is mostly due to the following ten centrality-dependent parameters $V,V_{1},V_{2},T,k,e_{2},n,f_{S},r_{xy},$ and $\rho_{0}$ which generally can depend also on the $p_{\textrm{T}}$, collision system and $\sqrt{s_{\mathrm{NN}}}$ ($f_{S}$ depends on particle species too). Typical number of the ALICE centrality classes for Pb–Pb data at $\sqrt{s_{\mathrm{NN}}}=$ 2.76 TeV [17] and 5.02 TeV [18] as well as for Xe–Xe data at $\sqrt{s_{\mathrm{NN}}}=$ 5.44 TeV  [19] is about 10 in each case. So, to describe these combined data, one needs to fit $10(10+10+10)=300$ independent values for the mentioned ten parameters. Taking into account also the RHIC data included in the present study, this number of independent values becomes more than 600. To reduce the number of free parameters strongly, we will parameterize these ten (and other) parameters by using the charged-particle multiplicities $M_{\rm ch}$ which depend on the collision system, centrality and $\sqrt{s_{\mathrm{NN}}}$, and are well known experimentally for various AA collisions  [17, 18, 19, 20, 21, 22]. The key assumption of the present paper is that all the centrality dependence in BWTPM can be expressed via the $M_{\rm ch}$ only. It is motivated by the experimental observation that any hadron species has similar $p_{\textrm{T}}$-spectra in different AA collisions with the same $M_{\rm ch}$ and $\sqrt{s_{\mathrm{NN}}}$ (see an example of such a similarity for ${\mathrm{K}}^{*0}$ meson $p_{\textrm{T}}$-spectra in Ref. [23]). To have a more general and predictive approach, we will use not the experimental values of $M_{\rm ch}$ (which can be missing for some centralities), but their accurate parameterization, valid for any AA collision centrality and energy. This parameterization is presented in the next section.
III parameterization of $M_{\rm ch}$
The mean value of charged-particle multiplicity density at midrapidity $M_{\rm ch}$ (defined in Eq. (2)) is one of the main characteristics of nucleus-nucleus collisions. Its dependence on the collision system, centrality and $\sqrt{s_{\mathrm{NN}}}$ was measured for different AA collisions at the LHC [17, 18, 19, 20] and RHIC [21, 22] energies. Centrality is commonly expressed in percentiles or fractions $x_{c}$ of the total nucleus-nucleus cross section [11] (centrality in % is equal to $100x_{c}$). Here we propose a simple parameterization for the $M_{\rm ch}$ data [17, 18, 19, 20, 21, 22] versus atomic mass number A, collision centrality fraction $x_{c}$ and energy $\sqrt{s_{\mathrm{NN}}}$. It relates $M_{\rm ch}$ with the similar quantity $M^{\rm pp}_{\rm ch}\equiv\langle\textrm{d}N^{\rm pp}_{\rm ch}/\textrm{d}\eta\rangle_{\eta=0}$ of the inelastic pp collisions at energy $\sqrt{s}=\sqrt{s_{\mathrm{NN}}}$. For $M^{\rm pp}_{\rm ch}$ we use the data measured at the LHC [24, 25, 26] as well as at lower energies [21, 27] and parameterize its energy dependence. The $M_{\rm ch}$ and $M^{\rm pp}_{\rm ch}$ data were obtained by averaging the respective multiplicity distribution on pseudo-rapidity in the range $|\eta|<0.5$ [17, 18, 19, 20, 22, 24, 25, 26] or $|\eta|<1$ [21, 27]. Within the data uncertainties both ranges give similar results [26]. Parameters of the parameterization are obtained via a simultaneous fit of the $M_{\rm ch}$ and $M^{\rm pp}_{\rm ch}$ data using the combined fit technique of the ROOT framework [28], which gives $\chi^{2}/NDF=54.1/155$. The fit results are shown in Figures 1 and 2 and the parameterization has the form
$\frac{M_{\rm ch}}{M^{\rm pp}_{\rm ch}}=1+(\textrm{A}-1)^{1+0.00013y}\frac{(1-x_{c}^{2})^{2+0.0018y\alpha}}{(x_{c}^{\delta}+\beta)^{2.1}\gamma},$ (13)
$M^{\rm pp}_{\rm ch}=0.615\frac{1+(\sqrt{s_{\mathrm{NN}}}/e_{0})^{0.258}}{1+(e_{0}/\sqrt{s_{\mathrm{NN}}})^{2}},$ (14)
where $\alpha=\log(\frac{17.4}{\sqrt{s_{\mathrm{NN}}}}),\beta=(\frac{0.0027}{\sqrt{s_{\mathrm{NN}}}})^{0.054},\delta=\textrm{A}^{0.063},\gamma=1+1.22x_{c}^{0.47},y=\textrm{A}(1-x_{c})$, and $e_{0}=0.005$ TeV (everywhere in the present paper, if it is not specified otherwise, $\sqrt{s_{\mathrm{NN}}}$ is defined in units of TeV). The bottom panel of Fig. 1 demonstrates that Eq. (13) describes most of the data-points with better than 10% accuracy. However, it underestimates the ALICE recent data for most peripheral Pb–Pb collisions at 5.36 TeV [20]. This could be due to the new centrality calibration [20], since all other ALICE data are described well. One can expect that Eq. (13) would give accurate values of $M_{\rm ch}$ also for new AA collisions. Result for oxygen–oxygen (OO) collisions, carried out at the LHC in July 2025, is also shown in Fig. 1. Note that for $\textrm{A}=1$ or for most peripheral collisions, when $x_{c}\to 1$, the right-hand side of Eq. (13) equals unity. For most central collisions, when $x_{c}\to 0$, we obtain a very simple formula for the maximal values of $M_{\rm ch}$ versus A and $\sqrt{s_{\mathrm{NN}}}$
$\frac{M_{\rm ch}^{max}}{M^{\rm pp}_{\rm ch}}=1+(\textrm{A}-1)^{1+0.00013\textrm{A}}(\frac{\sqrt{s_{\mathrm{NN}}}}{0.0027})^{0.1134}.$ (15)
For example, $M_{\rm ch}^{max}=$ 2404 (2457) for Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}=$5.02 (5.36) TeV.
Figure 1: Fit of the midrapidity charged-particle multiplicity density data measured for various AA collisions at different centralities $x_{c}$ and energies $\sqrt{s_{\mathrm{NN}}}$ by the ALICE [17, 18, 19, 20], PHOBOS [21] and PHENIX [22] experiments. Dashed line shows our prediction (multiplied by 0.5 for better visibility) for the OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV. Ratio data/fit demonstrates the quality of the fits.
To our knowledge, there is no similar to Eq. (13) generic parameterization in the literature. The existing parameterizations of $M_{\rm ch}$ versus $x_{c}$ (or versus the number of participant nucleons $N_{\mathrm{part}}$) are valid only for a particular AA collision system [21, 29]. Of course, there are many sophisticated models describing the $M_{\rm ch}$ data, most of which use the hydrodynamical approach for particle production in AA collisions (see, e.g., Ref. [30]). However, such models are much more complex with respect to our simple parameterization and their accuracy is worse.
Figure 2: Fit of the midrapidity charged-particle multiplicity density in inelastic pp collisions by Eq. (14) as a function of $\sqrt{s}$. Higher energy six data-points are from Refs. [24, 25, 26] and the remaining data-points are from Refs. [21, 27].
IV parameterization of BWTPM
Here we present parameterizations for the model parameters versus the $M_{\rm ch}$ (given by Eq. (13)), A, $\sqrt{s_{\mathrm{NN}}}$, and the particle mass and $p_{\textrm{T}}$. The parameterizations were chosen as simple as possible by examining the behavior of the model parameters versus the collision system, centrality, and energy, and using the results of Ref. [1] and other studies. For normalization parameters $V,V_{1}$, and $V_{2}$, to describe their strong increase with centrality [1] and taking into account the experimental fact that the hadron $p_{\textrm{T}}$-integrated yields increase linearly with $M_{\rm ch}$ at $M_{\rm ch}>100$ [31], the following parameterizations are assumed
$V=a_{1}M_{\rm ch}/(1+f_{r}M_{\rm ch}^{1/3}\frac{m}{m_{\textrm{T}}})^{a_{4}},~V_{1}=a_{2}M_{\rm ch},~V_{2}=a_{3}M_{\rm ch}.$ (16)
Here and throughout the present paper the parameters $a_{i},i=1,2,3,$... are fitting constants. Parameter $f_{r}$ is nonzero for ${\mathrm{K}}^{*0}$and $\rho^{0}$ only, and accounts for the suppression of the yields of these short-lived resonances at low $p_{\textrm{T}}$ due to the rescatterings of their decay products in the hadronic medium. The suppression is stronger for more central collisions and scales with $M_{\rm ch}^{1/3}$ [11].
The decrease of parameters $T$ and $k$ with increase of centrality and of parameters $e_{2}$ and $n$ with increase of centrality and $\sqrt{s_{\mathrm{NN}}}$ [1] can be parameterized as follows
$T=a_{5}/(1+a_{6}\log M_{\rm ch}),~k=a_{7}/(1+a_{8}M_{\rm ch}),$ (17)
$e_{2}=a_{9}/(1+a_{10}\log\sqrt{s_{\mathrm{NN}}})-a_{11}\log M_{\rm ch},$ (18)
$n=h_{n}[\frac{a_{12}}{1+a_{13}y\log\sqrt{s_{\mathrm{NN}}}}+\frac{1+(a_{14}/\sqrt{s_{\mathrm{NN}}})^{2y}}{(M_{\rm ch}/a_{15})^{a_{16}/(1+x^{2})}}],$ (19)
where $x=p_{\textrm{T}}/(a_{17}\log(10^{3}\sqrt{s_{\mathrm{NN}}})),y=(p_{\textrm{T}}/m_{\textrm{T}})^{a_{18}(c_{2}-1)}$ , $h_{n}$ differs from unity only for the strange and charmed baryons and $\mathrm{J}/\psi$ ($\sqrt{s_{\mathrm{NN}}}$ is in units of TeV). Parameter $c_{2}$ is given below in Eq. (27). The term with parameter $a_{14}$ is introduced to describe the RHIC data. It is significant at LHC energies for $p_{\textrm{T}}\ll m$ only. Note that $n$ becomes independent of $M_{\rm ch}$ at high $p_{\textrm{T}}$ where $x>1$.
The factor $f_{S}$ is equal unity for ordinary hadrons and increases with decreasing centrality (or $M_{\rm ch}$) for other hadrons. It is greater for hadrons containing larger number of strange and charm (anti)quarks, which we denote by $n_{s}$ and $n_{c}$, respectively. For example, $n_{s}=1$ for K and $\eta$  mesons, and $n_{s}=2$ for $\phi$ meson. We have found a simple parameterization for $f_{S}$ in AA collisions
$f_{S}=1+\frac{[a_{19}n_{s}+a_{20}n_{c}+a_{21}|B|(n_{s}+n_{c})]^{2}}{(1+p_{\textrm{T}}^{2}/a_{22}^{2})\sqrt{M_{\rm ch}+a_{23}}},$ (20)
which describes well the enhancement features of all considered hadrons containing at least one strange or charm (anti)quark ($n_{s}=n_{c}=0$ for ordinary hadrons). The term proportional to the hadron baryon number $B$ is introduced to better describe the production of hyperons. We assume the same parameter $a_{21}$ for strange and charmed baryons since the used data for the latter case are very limited. For the hadrons containing bottom quarks, not considered in the present study, the Eq. (20) can be modified accordingly. Note that $f_{S}$ tends to unity at $p_{\textrm{T}}>a_{22}$. Hence, the ratios of different particle yields in our model are independent of centrality at high $p_{\textrm{T}}$, as in the data [11].
It is known that the $p_{\textrm{T}}$-averaged $v_{2}$ increases when going from central to peripheral collisions, and then decreases for very peripheral collisions [11, 16]. Also, the eccentricity $\varepsilon_{2}$ of the transverse geometry of hadronic medium behaves similarly versus collision centrality. When going from the central to peripheral collisions, this geometry goes from circular to elliptical and $\varepsilon_{2}$ increases. But the geometry goes again to circular and $\varepsilon_{2}$ decreases at very peripheral collisions, where in average only one nucleon-nucleon collision accurs [10, 15]. So one can assume $v_{2}\sim\varepsilon_{2}$. This parameterization of $r_{xy}$ ensures such a centrality dependence for $\varepsilon_{2}$ (given by Eq. (7))
$r_{xy}=1-\frac{a_{24}\textrm{A}^{a_{25}}}{1+a_{26}/x+a_{27}x^{a_{29}}+a_{28}[(1-f_{d})x]^{a_{30}}},$ (21)
where $x=\textrm{A}^{a_{25}}(z_{c}+y)/(1+y)$, A is the atomic mass number, $z_{c}=M_{\rm ch}/M_{\rm ch}^{0\textrm{-}5\%}$ is the normalized multiplicity over the multiplicity of the 0-5% most central collisions $M_{\rm ch}^{0\textrm{-}5\%}$ (determined by Eq. (13) at $x_{c}=0.025$) and its correction $y=(a_{31}/\sqrt{s_{\mathrm{NN}}})^{a_{32}}z_{c}$ is needed mainly for the fit of $v_{2}$ data at RHIC energies. Parameter $f_{d}$ is introduced to describe a slightly different centrality dependence of $v_{2}$ data in Xe–Xe collisions compared to all other collision systems considered in present study for which we use $f_{d}=0$. The difference is related to the non-spherical (deformed) shape of 129Xe nucleus [32].
To describe the centrality and energy dependence of the hadrons $p_{\textrm{T}}$-spectra and also of the charged-particle average transverse momentum $\langle p_{\textrm{T}}\rangle$ [33, 34], particularly its local minimum at $z_{c}\sim 1.05$ and rise at $z_{c}>1.05$ for ultracentral collisions [34], we parameterize $\rho_{0}$ as
$\rho_{0}=\frac{1}{f_{uc}}\bigl{[}\frac{a_{33}\sqrt{s_{\mathrm{NN}}}^{a_{34}}}{1+f_{d}\frac{M_{\rm ch}}{a_{35}}}-a_{36}\log^{2}(\frac{z_{c}}{a_{37}})\bigr{]},$
$f_{uc}=\biggl{[}1-\frac{a_{38}}{z_{c}}\sqrt{\frac{2}{\pi}}\frac{\exp(-\frac{(z_{c}-a_{39})^{2}}{2a_{38}^{2}})}{\textrm{erfc}(\frac{z_{c}-a_{39}}{\sqrt{2}a_{38}})}\biggr{]}^{a_{40}}.$ (22)
Factor $f_{uc}$ [34, 35] is close to unity at $z_{c}\!\leq\!a_{38}\!\sim\!1.1$ and decreases at $z_{c}\!>\!a_{39}$. So, it increases $\rho_{0}$ at $z_{c}\!>\!a_{39}$ and therefore increases $\langle p_{\textrm{T}}\rangle$ for ultracentral events like in Refs. [34, 35]. Parameter $a_{35}$ is used to describe the observed slightly different centrality dependence of $\langle p_{\textrm{T}}\rangle$ in Xe–Xe and Pb–Pb collisions which is due to the non-sphericity of 129Xe nuclei [33].
So, the parameterizations in Eqs. (16)–(IV) allow us to determine the aforementioned in Sec. II more than 600 independent values for $V,V_{1},V_{2},T,k,e_{2},n,f_{S},r_{xy},$ and $\rho_{0}$ by means of about 40 constants.
We further parameterize also other ingredients of BWTPM. The energy dependence of parameter $f_{hyp}$ from Eqs. (11) and (12), and of the global normalization $f_{N}$ are given as
$f_{N}=a_{N}[1+(a_{41}/\sqrt{s_{\mathrm{NN}}})^{a_{42}}],~f_{hyp}=\sqrt{s_{\mathrm{NN}}}/a_{43},$ (23)
where $a_{N}=1$ is assumed for Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}=$ 5.02 TeV and for other used data-sets the values of $a_{N}$ are defined relative to this one. As was mentioned in Sec. II, we assume that the parameter $\rho_{2}$ as well as the functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ are proportional to $\varepsilon_{2}$ in order to ensure the relation $v_{2}\sim\varepsilon_{2}$ [13, 14, 15, 16]. A weak energy dependence of $\rho_{2}$ is parameterized as
$\rho_{2}=\varepsilon_{2}a_{44}/\sqrt{s_{\mathrm{NN}}}^{a_{45}},$ (24)
which becomes significant at RHIC energies.
Functions $v_{2t}(p_{\textrm{T}})$ and $v_{2p}(p_{\textrm{T}})$ are responsible mainly for the description of the high-$p_{\textrm{T}}$ behavior of elliptic flow $v_{2}$. For their dependence on $p_{\textrm{T}}$, $\sqrt{s_{\mathrm{NN}}}$, and $M_{\rm ch}$ we use
$v_{2t}(p_{\textrm{T}})=\varepsilon_{2}\frac{a_{46}}{y}\frac{x^{z}}{1+t^{2}},~~v_{2p}(p_{\textrm{T}})=\varepsilon_{2}\frac{a_{47}}{y}\frac{x^{z}}{1+t^{2}},$ (25)
[…]
$\thinspace\frac{p_{2}}{\textrm{GeV}/c}\!=\!\left\{\begin{array}[]{ll}a_{51}+a_{52}(1-\frac{m_{\pi}}{m})\log(\frac{a_{53}}{m})\;\textrm{for mesons},\\ a_{54}+a_{55}\exp(\frac{a_{56}}{m})\qquad\qquad\textrm{for baryons},\!\!\\ a_{57}\qquad\qquad\qquad\qquad\qquad\;\;\textrm{for }\mathrm{J}/\psi.\end{array}\right$. (26)
parameterization for mesons is inaccurate for $\mathrm{J}/\psi$.
The particle type dependence of parameter $c_{2}$ can be expressed by a simple relation
$c_{2}=1+\frac{m}{\textrm{GeV}/c^{2}}\left\{\begin{array}[]{ll}a_{58}H(m-m_{\pi})\quad\textrm{for mesons},\\ a_{59}H(m-m_{\mathrm{p}})\quad\textrm{for baryons},\end{array}\right$. (27)
where $m_{\mathrm{p}}$ is the proton mass and the step function $H$ is defined as $H(x)=0(1)$ for $x\leq(>)~0$. So $c_{2}=1$ for pions and (anti)proton. Larger $a_{59}$ with respect to $a_{58}$ is nedded to describe the baryon yields. Note that the different parameterizations for mesons and baryons in Eqs.  (26) and (27) indicate the difference in production mechanisms of these hadrons containing two and three (anti)quarks, respectively.
Parameter $e_{1}$ shows a centrality dependence significant mainly for pions. It can be described by
$e_{1}=\frac{\tilde{e}_{1}}{1+\sqrt{f_{\mathrm{p}}/\sqrt{s_{\mathrm{NN}}}}}(1+\frac{m_{\pi}}{m}\frac{a_{60}}{M_{\rm ch}+a_{23}}),$ (28)
where parameter $\tilde{e}_{1}$ depends on the particle species. Parameter $f_{\mathrm{p}}$ is nonzero for (anti)proton only to better describe its $p_{\textrm{T}}$-spectra at RHIC energies. Additional normalization parameter $f_{\rm D}$ is used for $\mathrm{D}^{\pm}$ mesons to describe their lower yield with respect to ${\mathrm{D}}^{0}$ meson.
V Fitted data and parameters
Thus, we presented in previous Sections all the ingredients of the BWTPM. To determine the values of its free parameters, a global simultaneous fit is performed of the combined data on various hadrons midrapidity $p_{\textrm{T}}$-spectra, $\langle p_{\textrm{T}}\rangle$, and $v_{2}$ for different centrality classes in Pb–Pb collisions at energies $\sqrt{s_{\mathrm{NN}}}=2.76$ and 5.02 TeV [18, 33, 34, 36, 37, 40, 41, 42, 38, 39, 43, 44, 45, 48, 49, 46, 47, 50, 51, 54, 55, 52, 53, 56, 57, 58, 59, 60, 61] and in Xe–Xe collisions at 5.44 TeV [32, 62, 63, 64]. To check the validity of BWTPM for other colliding nuclei and at much lower energies, we included in this fit also the $p_{\textrm{T}}$-spectra and $v_{2}$ for $\pi^{\pm}$, $\mathrm{K}^{\pm}$ and $\mathrm{p}$($\mathrm{\overline{p}}$) measured in Au–Au collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV [65, 66], and the $p_{\textrm{T}}$-spectra for $\pi^{\pm}$ and $\mathrm{K}^{\pm}$ measured in Au–Au collisions at 39 and 27 GeV [67], in UU collisions at 193 GeV [68] and in Cu–Cu collisions at 200 GeV [69]. In the last four cases the proton data not included since they are not corrected for the contribution from weak decays of hyperons. The resulting ratio $\chi^{2}/NDF$ for this global fit (in the ROOT framework [28]) is 6738/13408. Values for the fitting parameters are given in Tables I–III. Parameters $f_{BW},f_{1}$ and $f_{2}$ for $\pi^{\pm}$ and $\pi^{0}$ are fixed equal to unity. Note that in the fits we use the data of prompt $\rm D$} and $\mathrm{J}/\psi$ mesons and $\Lambda_{c}$ baryons, not including contributions from decays of heavier hadrons, containing $b$-quarks. To describe the $p_{\textrm{T}}$-spectra in Xe–Xe collisions [62, 63], we scale the right-hand side of Eq. (12) by a factor of 1.07 to account for a small normalization difference between the measured $p_{\textrm{T}}$-spectra of identified [62] and unidentified charged particles [63]. Note also that parameters $a_{24}$$a_{32}$ and $a_{44}$$a_{57}$ are resposible mainly for the elliptic flow. Only some of them entering in the Eq. (24) influence the azimuthally integrated $p_{\textrm{T}}$-spectra due to the $\rho_{2}$ contribution into the radial flow velocity Eq. (5). This contribution is small and decreases with increasing energy, according to Eq. (24). So, its influence on $p_{\textrm{T}}$-spectra is negligible at high energies and becomes sizable (at most about several % at high $p_{\textrm{T}}$) only for Au–Au collisions at low energies $\sqrt{s_{\mathrm{NN}}}=27$ and 39 GeV.
Table 1: Particle-type dependent model parameters.
parameter $\pi^{\pm}$, $\pi^{0}$ $\mathrm{K}^{\pm}$, ${\mathrm{K}}^{0}_{\mathrm{S}}$ $\mathrm{p}$, $\mathrm{\overline{p}}$ $\eta$ $\rho^{0}$ ${\mathrm{K}}^{*0}$, $\overline{\mathrm{K}^{*0}\mkern-15.0mu}$ $\phi(1020)$ $\Lambda$, $\overline{\Lambda}$ $\Xi^{-}$, $\overline{\Xi^{+}\mkern-12.0mu}$ $\Omega^{-}$, $\overline{\Omega^{+}\mkern-12.0mu}$ ${\mathrm{D}}^{0}$, $\mathrm{D}^{\pm}$ $\mathrm{J}/\psi$ $\Lambda_{c}^{+}$$\overline{\Lambda^{-}\mkern-12.0mu}_{c}$
$f_{BW}$ 1 0.993 3.76 1.26 2.18 2.23 1.66 10.96 7.08 8.29 192 1500 3600
$f_{1}$ 1 0.085 0.00015 0 0 0 0 0 0 0 0 0 0
$f_{2}$ 1 0.528 0.041 0.490 0.360 0.163 0.117 4.70 1.64 0.428 0.735 0.111 6.0
$c_{1}$ 1.30 1.43 3.06 - - - - - - - - - -
$\tilde{e}_{1}$ [GeV] 0.065 0.099 0.361 - - - - - - - - - -
$h_{n}$ 1 1 1 1 1 1 1 1.30 1.30 1.30 1 1.14 1.20
Table 2: Parameters $a_{1}$$a_{60}$ of Eqs. (16)–(28), $f_{r}$ (is nonzero for $\rho^{0}$ and ${\mathrm{K}}^{*0}$only) of Eq. (16) , $f_{d}$ (is nonzero for Xe–Xe only) of Eqs. (21) and  (IV), $f_{\mathrm{p}}$ (is nonzero for $\mathrm{p}$ and $\mathrm{\overline{p}}$ only) of Eq. (28), and $f_{\rm D}$ for normalization of $\mathrm{D}^{\pm}$ mesons.
$a_{1}[{\rm fm}^{3}]$ $a_{2}[{\rm fm}^{3}]$ $a_{3}[{\rm fm}^{3}]$ $a_{4}$ $a_{5}$ [GeV] $a_{6}$ $a_{7}$ $a_{8}$ $a_{9}$ [GeV] $a_{10}$ $a_{11}$ [GeV] $a_{12}$ $a_{13}$
6.15 17.62 0.015 2.0 0.1354 0.020 2.61 0.012 1.477 0.015 0.059 7.66 0.080
$a_{14}$ [TeV] $a_{15}$ $a_{16}$ $a_{17}$ [GeV] $a_{18}$ $a_{19}$ $a_{20}$ $a_{21}$ $a_{22}$ [GeV] $a_{23}$ $a_{24}$ $a_{25}$ $a_{26}$
0.027 10.85 0.80 35.0 3.60 0.92 2.50 0.53 24.0 5.0 0.341 0.070 0.0042
$a_{27}$ $a_{28}$ $a_{29}$ $a_{30}$ $a_{31}$ [TeV] $a_{32}$ $a_{33}$ $a_{34}$ $a_{35}$ $a_{36}$ $a_{37}$ $a_{38}$ $a_{39}$
0.782 0.159 1.56 6.63 0.166 1.80 0.984 0.083 5100 0.014 0.81 0.014 1.103
$a_{40}$ $a_{41}$ [TeV] $a_{42}$ $a_{43}$ [TeV] $a_{44}$ $a_{45}$ $a_{46}$ $a_{47}$ $a_{48}$ $a_{49}$ $a_{50}$ $a_{51}$ $a_{52}$
0.301 0.02 0.52 21.5 0.386 0.131 6.13 2.13 0.166 0.689 0.042 3.04 1.14
$a_{53}$ [GeV] $a_{54}$ $a_{55}$ $a_{56}$ [GeV] $a_{57}$ $a_{58}$ $a_{59}$ $a_{60}$ $f_{r}$, $\rho^{0}$ $f_{r}$, ${\mathrm{K}}^{*0}$ $f_{d}$, Xe–Xe $f_{\mathrm{p}}$ [TeV] $f_{\rm D}$
0.76 1.60 0.05 3.63 2.67 0.305 0.595 1.70 0.112 0.112 0.133 0.009 0.45
Table 3: Normalization parameter $a_{N}$ defined in Eq. (23) for all considered collision systems. Corresponding $\sqrt{s_{\mathrm{NN}}}$ values are given in units of TeV.
Pb–Pb Xe–Xe Au–Au $U$ Cu–Cu
$\sqrt{s_{\mathrm{NN}}}$ 5.02 2.76 5.44 0.2 0.039 0.027 0.193 0.2
$a_{N}$ 1 1.044 0.986 1.008 1.003 0.973 1.110 0.965
VI Discussion of the results
VI.1 Hadrons $p_{\textrm{T}}$-spectra, integrated yields, and $\langle p_{\textrm{T}}\rangle$
First, we present several plots to illustrate the results and quality of the $p_{\textrm{T}}$-spectra fits. Figure 3 displays the fits of different data sets with highest $p_{\textrm{T}}$ reach in central Pb–Pb collisions. To demonstrate the quality of the fits, the data points have been divided by the corresponding fit function values, and these ratios are also plotted in the bottom panel. Generally, the quality is good within the data uncertainties. Figures 4, 5, and 6 display our main fits of the centrality dependent spectra for most abundantly produced pions, kaons, and protons. The fit quality is almost always very good. This is true also for the fits of $p_{\textrm{T}}$-spectra of all considered other particle species. Our predictions for the 0-5% most central OO collisions are shown in Fig. 5. Note that Figs. 3 and 4
Figure 3: Fit of $p_{\textrm{T}}$-spectra with highest $p_{\textrm{T}}$ reach in Pb–Pb collisions for charged particles at $\sqrt{s_{\mathrm{NN}}}$ ==2.76 TeV [37, 40] and 5.02 TeV [37, 41] and for ${\mathrm{D}}^{0}$  at $\sqrt{s_{\mathrm{NN}}}$ ==5.02 TeV [42, 47].
Figure 4: Fit of pion, kaon, and proton $p_{\textrm{T}}$-spectra at $|y|<$ 0.5 for different centrality classes (centrality intervals) in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV [18]. The data points and fitting curves in the top panels are scaled by the numbers given in the parentheses for a better visibility.
Figure 5: Similar to the Fig. 4 but for Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [62]. In addition, the dashed lines in top panels show our predictions (scaled down by 0.0002 for better visibility) for 0-5 % centrality OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
Figure 6: Similar to the Fig. 4 but for Au–Au collisions at $\sqrt{s_{\mathrm{NN}}}$ == 0.2 TeV [65].
repeat the similar ones of Ref. [1] to show that the present version of BWTPM with much less number of fitting parameters gives similar fit quality. Also, we have checked that the relative contributions of the BW, Ts, and Po components of present model into the $p_{\textrm{T}}$-spectra are similar to the ones given in Ref. [1]. As shown in Table I, the contribution of the Ts-component is nonzero only for pions, kaons, and protons, which have significant feedback from the resonance decays.
Then, we show in Fig. 7 a good agreement of the model calculations (using the parameter values given in Sec. V) for the ratios of $p_{\textrm{T}}$-spectra of different hadrons with the ALICE data obtained in central and peripheral Pb–Pb collisions. These ratios, like the other ones given in Ref. [1], demonstrate more clearly the effect of hadron production mechanisms on the $p_{\textrm{T}}$-spectra. The peaks at $p_{\textrm{T}}$ $\sim$ 2–7 GeV/c are related to the radial flow, which is stronger for heavier particles and more central collisions. The ratios with the short-lived ${\mathrm{K}}^{*0}$and $\rho^{0}$ resonances show the suppression of their yields for low $p_{\textrm{T}}$ and central collisions, according to Eq. (16). To demonstrate the collision system and energy dependence of particle ratios, we show in the left panels of Fig. 7 the model calculations for 0-10% centrality Au–Au (dashed lines) and Cu–Cu (dotted lines) collisions at $\sqrt{s_{\mathrm{NN}}}$ == 200 GeV. Both dependencies are not strong and appear at moderate $p_{\textrm{T}}$. Note that the hadron $p_{\textrm{T}}$-spectra in Eq. (II) have a power-law behavior $\propto p_{\textrm{T}}^{2-n}$ at high $p_{\textrm{T}}$. Hence, the ratios for all hadrons, which have $h_{n}=1$ in Table I, reach a plateau (independent of the collision system, centrality, and $\sqrt{s_{\mathrm{NN}}}$) at high $p_{\textrm{T}}$ where parameter $n$ in Eq. (19) depends only on $\sqrt{s_{\mathrm{NN}}}$. The plateau is simply defined by the ratio of factors $(2J+1)f_{2}$ with $f_{2}$ given in Table I. It is true also for the ratios between hyperons which have the same $h_{n}$. For example, the ratio $\eta$$\pi^{0}$ reach a plateau value of 0.49, which is compatible with the measurement $0.457\pm 0.013\pm 0.018$ [57] and with the empirical fit result $0.487\pm 0.024$ [70].
Figure 8 displays the ratios of the $p_{\textrm{T}}$-integrated yields of various hadrons as a function of the charged-particle multiplicity $M_{\rm ch}$ (defined by Eq. (2)) in Pb–Pb collisions. It is well known that such ratios have smooth behavior versus $M_{\rm ch}$ and are almost independent of the collision system and energy [11, 32]. Corresponding ratios in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [32] are not shown in Fig. 8 for its clearity. Our model curves in this case, as well as for OO collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.36 TeV, coincide practically with the ones for Pb–Pb at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV. The model describes well all the presented ratios. In particular, it describes the observed enhancement of yields of hadrons containing strange (anti)quarks with respect to pion yields when going from peripheral to central collisions. Enhancement is stronger for hadrons containing more strange (anti)quarks, according to Eq. (20). Figure 8 displays also the suppression of ${\mathrm{K}}^{*0}$and $\rho^{0}$ resonance yields when going from peripheral to central collisions. It is the manifestation of the low-$p_{\textrm{T}}$ suppression of these resonances (see Fig. 7) due to the rescatterings of their decay products in the hadronic medium. The $\phi$ resonance does not have such a suppression since it lives longer than this medium [11].
Next, we show in Fig. 9 the fit of the data on charged-particle $\langle p_{\textrm{T}}\rangle$ versus $M_{\rm ch}$ in Pb–Pb and Xe–Xe collisions [33]. These data involve much more values of $M_{\rm ch}$
Figure 7: Comparison of the model calculations for the ratios of different particle $p_{\textrm{T}}$-spectra with the data of central and peripheral Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 and 5.02 TeV [39, 57, 58, 59]. The dashed (dotted) lines show the ratios for 0-10% centrality Au–Au (Cu–Cu) collisions at $\sqrt{s_{\mathrm{NN}}}$ == 200 GeV.
Figure 8: Comparison of the model calculations for the ratios of different particle $p_{\textrm{T}}$-integrated yields as a function of the charged-particle multiplicity density with the measurements in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 and 5.02 TeV [17, 18, 39, 43, 44, 45].
Figure 9: Fit of charged particles $\langle p_{\textrm{T}}\rangle$ as a function of their multiplicity density in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 and 5.02 TeV, and in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [33]. The dashed line shows our prediction for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
Figure 10: Fit of charged particles normalized $\langle p_{\textrm{T}}\rangle$ as a function of their normalized multiplicity density $z_{c}$ in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV [34]. The predictions for Pb–Pb at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 TeV, Xe–Xe at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV, and OO at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV are shown too.
than the $p_{\textrm{T}}$-spectra measurements [37, 63]. Our prediction for OO collisions is shown too. The sharp rise in model curves at very high $M_{\rm ch}$, related to factor $f_{uc}$ of Eq. (IV), is demonstrated in detail in Fig. 10 with the Pb–Pb data [34] at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV for the normalaized $\langle p_{\textrm{T}}\rangle$ and $M_{\rm ch}$ to the corresponding values for 0-5% centrality. Our predictions for Pb–Pb at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 TeV, Xe–Xe at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV, and OO at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV are also shown. Overall, a good description of the data can be seen.
VI.2 Hadrons $p_{\textrm{T}}$-differential and averaged $v_{2}$
It should be noted that the elliptic flow $v_{2}$ data considered in present paper [32, 51, 52, 53, 54, 55, 56, 60, 61, 64, 67] were measured mostly by the two-particle cumulant method and partly by other close methods that give compatible results within the data uncertainties. This method assumes a minimum pseudo-rapidity gap $\Delta\eta$ between the two particles, meant to suppress the non-flow correlations, for which we have chosen the largest available value in any dataset used.
Figure 11: Fit of pion, kaon, proton, and charged-particle $v_{2}$ as a function of $p_{\textrm{T}}$ in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 TeV for four centrality intervals, measured by ALICE [52, 53] and CMS [54] experiments. The dashed lines show our predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
Figures 11, 12, and 13 demonstrate our fits of $p_{\textrm{T}}$-differential $v_{2}$ for various hadrons and different centralities in Pb–Pb [51, 52, 53, 54] and Au–Au [67] collisions, as well as some features of such data [11, 60]. A similar good fit is also obtained for the $v_{2}$ data in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV [32, 64]. The $v_{2}$ ($p_{\textrm{T}}$) has a wide peak at $p_{\textrm{T}}\sim$ 1–7 GeV/c which height is increasing with the percentage of centrality up to $\sim$50% and then decreasing for more peripheral collisions. This is related to the collision geometry [60] and can be explained in present model by the similar centrality dependence of $\varepsilon_{2}$, defined by Eqs. (7) and (21). The peak is shifted to higher $p_{\textrm{T}}$ for heavier particles. It is the well-known mass-ordering effect due to the radial flow [11, 60]. Such a shift also occurs with increasing centrality (or $M_{\rm ch}$) due to the increasing radial flow velocity $\beta_{\textrm{T}}$ in Eq. (5) caused by the decreasing exponent $k$ defined in Eq. (17). Other known feature of $v_{2}$ is the meson-baryon particle type grouping [11, 60], clearly visible in Fig. 12. This is described in present model by using different parameterizations for mesons and baryons in Eq. (26). Another effect is the increase of $v_{2}$ with $\sqrt{s_{\mathrm{NN}}}$ [11] which can be seen by comparing Figs. 11 and 13. It is mainly related to the increase of $\beta_{\textrm{T}}$ caused by the decreasing $k$ in Eq. (17) with increasing $\sqrt{s_{\mathrm{NN}}}$ (or $M_{\rm ch}$). For the fit shown in Fig. 13 we have used the $\pi^{-}$, $\mathrm{K}^{-}$, and $\mathrm{\overline{p}}$ dataset since it includes more
Figure 12: Fit of $v_{2}$ for different particles as a function of $p_{\textrm{T}}$ in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 2.76 TeV for centrality intervals 10-20% and 40-50% [51].
Figure 13: Fit of $v_{2}$ for $\pi^{-}$, $\mathrm{K}^{-}$, and $\mathrm{\overline{p}}$  as a function of $p_{\textrm{T}}$ in Au–Au collisions at $\sqrt{s_{\mathrm{NN}}}$ == 200 GeV for three centrality intervals [67].
data-points than the $\pi^{+}$, $\mathrm{K}^{+}$, and $\mathrm{p}$ dataset [67]. Note also that the dashed lines in Fig. 11 show our predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
It is important to note that the BW term of our model (see Eq. (II)) gives, for the most of considered hadrons, the main contribution to $v_{2}$ in the lower part of its whole $p_{\textrm{T}}$-region up to the center of the peak. The Ts and Po terms, with the use of Eq. (25), describe $v_{2}$ in the rest of this region. However, for hadrons containing heavy quarks, for which the Ts term is zero and the BW term is suppressed by a factor $\sim\exp(-m_{\textrm{T}}/T)$, the Po contribution into $v_{2}$ dominates everywhere (like for $p_{\textrm{T}}$-spectra [1]).
Finally, let us discuss the $p_{\textrm{T}}$-averaged $v_{2}$, defined via Eq. (1) with integration by $p_{\textrm{T}}$ in the numerator and denominator. Figure 14 shows our fit for the charged-particle $v_{2}$, averaged in the $p_{\textrm{T}}$ range $0.2<p_{\textrm{T}}<3.0$ GeV/c, as a function of $M_{\rm ch}$ [61] (left panel) and centrality fraction $x_{c}$ (defined in Sec. III[53] (right panel). Although both datasets were obtained by the two-particle cumulant method, they are independent measurements since Refs. [61] and [53] use $\Delta\eta>1.4$ and $\Delta\eta>1$, respectively. Very good fits of the Pb–Pb and Xe–Xe data can be seen in Fig. 14. The black lines show our predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV.
Figure 14: Fit of charged-particle $v_{2}$, averaged in the $p_{\textrm{T}}$ range $0.2<p_{\textrm{T}}<3.0$ GeV/c, as a function of multiplicity density [61] (left panel) and $x_{c}$ [53] (right panel) in Pb–Pb collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.02 and 2.76 TeV, and in Xe–Xe collisions at $\sqrt{s_{\mathrm{NN}}}$ == 5.44 TeV. The black lines show our predictions for OO collisions at $\sqrt{s_{\mathrm{NN}}}=$5.36 TeV. The dashed lines correspond to $v_{2}$ averaged in the $p_{\textrm{T}}$ range $0<p_{\textrm{T}}<20$ GeV/c.
For comparison, dashed lines are shown, corresponding to $v_{2}$ averaged in the much larger $p_{\textrm{T}}$ range $0<p_{\textrm{T}}<20$ GeV/c. The left panel of Fig. 14 shows that our model describes well the different $M_{\rm ch}$ (or centrality) dependence of $v_{2}$ in Xe–Xe and Pb–Pb collisions. This is due to the presence of the parameter $f_{d}$, related to the non-spherical form of 129Xe nucleus, and of the A-dependence in Eq. (21). Also important is the smaller $\beta_{\textrm{T}}$ (caused by the larger $k$ from Eq. (17)) for collisions of smaller nuclei. These A-dependence and smaller $\beta_{\textrm{T}}$ are the main reasons for the lower values and flatter centrality (or $M_{\rm ch}$) dependence of the predicted $v_{2}$ in OO collisions with respect to Pb–Pb collisions, shown in Figs. 11 and 14.
VII Conclusions
Thus, a more general and parameterized version of BWTPM [1] is presented, which has much fewer free parameters, but successfully describes not only the midrapidity $p_{\textrm{T}}$-spectra, but also elliptic flow $v_{2}$ of various hadrons with arbitrary $p_{\textrm{T}}$ in AA collisions for any centrality at the LHC energies. It is shown that the model also works at much lower RHIC energies. For all considered particles, from pions to charmonia, BWTPM allows us to calculate easily the midrapidity $p_{\textrm{T}}$-spectra and $v_{2}$ for new AA collisions with any A and centrality at high energies. Predictions are given for the OO collisions carried out recently at the LHC. To calculate the $p_{\textrm{T}}$-spectra or $v_{2}$ for new particles, not considered here, one needs some data (e.g., $p_{\textrm{T}}$-spectrum for one centrality or $\langle p_{\textrm{T}}\rangle$ versus centrality) to fix at least two particle-type dependent parameters $f_{BW}$ and $f_{2}$.
BWTPM can be efficiently used by experimenters to quickly check new data on hadron $p_{\textrm{T}}$-spectra and $v_{2}$, and by theorists to test their more general and sophisticated theories and models of hadron production in AA collisions. Similarly, the simple parameterization for the charged-particle multiplicity density at midrapidity in AA or pp collisions presented in Sec. III can also be used.
Acknowledgments
I thank Jurgen Schukraft for valuable comments on the BWTPM parameterization.
Data Availability
The data supporting the findings of this article are publicly available and cited in the bibliography.